# Lifelong Learning with Behavior Consolidation for Vehicle Routing

### v2

@@ 3.3 Decision-seeking Behavior Consolidation (DsBC)
By minimizing the weighted sum of the difference between the current model’s behaviors and the buffered behavior at previously encountered states, LLR-BC encourages the consolidation of prior behavior patterns, thereby preserving knowledge and mitigating catastrophic forgetting.
The Kullback–Leibler divergence (KLD) is widely adopted for knowledge distillation of neural VRP solvers (Bi et al., 2022; Zheng et al., 2025). However, recent studies (Kaplanis et al., 2019; Wang et al., 2024a; Qi et al., 2025) have shown that using reverse KLD (RKLD) could lead to better learning performance. Minimizing KLD pushes the learner distribution to spread probability mass across all modes of the given teacher distribution, i.e., mode-covering (Wang et al., 2024a; Qi et al., 2025). In contrast, minimizing RKLD leads to mode-seeking, where the learner tends to concentrate on the teacher’s highest-probability actions, while still maintaining relatively good alignment with the overall distribution (Wang et al., 2024a; Qi et al., 2025). During problem-solving of constructive neural VRP solvers, the action (selected node) with the highest probability is typically chosen to extend the route. To prevent catastrophic forgetting, that could manifest in significantly longer routes on previously seen tasks, it is crucial to preserve these highest-probability decisions. Therefore, instead of KLD, LLR-BC employs the RKLD to measure the behavioral difference in a decision-seeking way. To consolidate sampled behaviors, LLR-BC minimizes the following behavior consolidation loss term, considering the RKLD (Wang et al., 2024a; Qi et al., 2025):
@@
where $\{\tau\}$ is the set of newly obtained experience trajectories from the new task, $\mathcal{L}_{DRL}(\theta,\{\tau\})$ is the loss calculated by the adopted DRL algorithm based on $\{\tau\}$ to maximize the reward gain, and $\alpha$ is a hyperparameter to balance between consolidation of previous experiences and learning of the new task. Notably, LLR-BC is a generic framework and can be applied to different neural solver methods with different DRL algorithms, in which $\mathcal{L}_{DRL}$ in Eq. equation 2 can be implemented in different ways. A detailed discussion about computational cost of DsBC can be found in Appendix A.4.
@@ 4 Experiments
  • Stability and plasticity: Detail in lifelong learning process, how does LLR-BC perform in terms of stability and plasticity?
@@
To answer the questions, we simulate multiple lifelong learning scenarios on CVRP and TSP, and compare LLR-BC against widely used and representative lifelong learning baselines across diverse metrics. Most experiments are conducted based on POMO (Kwon et al., 2020), given its concise design and broad applicability. To verify the applicability of LLR-BC, we further evaluate LLR-BC on two representative and state-of-the-art constructive neural solvers designed for good cross-distribution cross-scale performance, i.e., Omni (Zhou et al., 2023) and INViT (Fang et al., 2024).
@@
Six distributions for simpling node coordinates are used: four proposed by  Bossek et al. (2019) and widely used in neural VRP solver stuides (Jiang et al., 2022; Zhou et al., 2023), i.e., Uniform (U), Gaussian Mixture (GM), Explosion (E), Compression (C), and two additionally designed ones for greater diversity: Grid (G) and Ring (R). For CVRP, node demand is also required to be sampled. While prior studies used uniformly random demand for all distributions (Kwon et al., 2020; Jiang et al., 2022; Zhou et al., 2023; Fang et al., 2024), we build six distinct demand distributions and assign them to the above six distributions to better reflect the diversity of real-world problems (Liu et al., 2021). Following common practice (Kwon et al., 2020; Jiang et al., 2022), we consider three problem scales: 20, 50, and 100. Specifically, tasks U and R correspond to scale 20, tasks G and E to scale 50, and tasks C and GM to scale 100, arbitrarily. We construct five task orders by randomly permuting the tasks. In addition, two classic and widely used benchmark datasets, TSPLIB (Reinelt, 1991) and CVRPLIB (Uchoa et al., 2017), are involved for generalization ability evaluation. Appendix B provides the full definition of tasks and orders.
@@
Following Kwon et al. (2020), we train the solver on each task over 200 epochs, solve a batch of sampled problems in parallel to leverage GPU parallelism during training. Therefore, both experience buffering and sampling operate at the batch level rather than the single experience level. Hyperparameters of LLR-BC are set as follows: $|\mathcal{B}|=1000$, $|\mathcal{E}|=16$, and $\alpha=100$, where $|\mathcal{B}|$ and $|\mathcal{E}|$ are defined in units of experience batches. The buffer constitutes only about 0.01% of the total training experiences obtained across all tasks. For evaluation, each task has a test set of 1,000 problem instances, which are different from the training ones. For each trial, all methods are evaluated on the same test sets. Additional details are provided in Appendix C.
@@
The goal of this work is not to design a totally new solver that outperforms the state-of-the-art solver on each individual task or a fixed set of tasks that can be learned simultaneously, but to improve the solver’s ability to learn multiple unpredictable tasks continuously. To this end, we consider several representative strategies, including two commonly adopted strategies in neural solver studies (Kwon et al., 2020; Zhou et al., 2023): (i) Restart, reinitialize the solver and train from scratch on each task; (ii) Fine-tuning, simply sequential training on each task. Additionally, we evaluate (iii) EWC (Kirkpatrick et al., 2017), a widely used method in lifelong learning, especially in recent studies of optimization tasks (Manchanda & Ranu, 2023; Pei et al., 2025a). We adapt and compare with (iv) LiBOG (Pei et al., 2025a), a recent method originally proposed for black-box optimization, within the VRP tasks. Appendix C.1 provides more details.
@@
  • Average Max Forgetting (AMF), the average of maximal forgetting of each task during the lifelong learning, i.e., $AMF=\frac{1}{k-1}\sum_{i=1}^{k-1}\max_{j=i+1}^{k}\max(0,\bar{d}_{j,i}-\bar{d}_{i,i})$.
@@
Smaller values of above metrics indicate better performance. Unless otherwise specified, all reported metric values in tables are scaled by $10^{-3}$.
@@ 4.1 Performance in Solving Seen Tasks
The experimental results for CVRP and TSP are summarized in Table 1, with $k=6$ for all metrics. More results can be found in Appendices D.1, D.2. Across all task orders, LLR-BC achieves average AP values of $0.0042$ (CVRP) and $0.0034$ (TSP), whereas all other methods exceed $0.023$ (CVRP) and $0.014$ (TSP), respectively, indicating that LLR-BC obtains solutions of problems from learned tasks with substantially better quality. Moreover, LLR-BC exhibits greater robustness to task order variation. Compared methods show large fluctuations (larger Std.) in AP across orders, while LLR-BC consistently maintains low AP values with lower standard deviations. These results demonstrate that LLR-BC effectively learns from sequentially arising tasks and retains the ability to solve all encountered tasks with high effectiveness.
@@
Method CVRP TSP AP AF AMF APl AG AP AF AMF APl AG Restart 60.5 (29.3) 41.3 (26.3) 52.0 (25.5) 9.1 (-) 49.5 (4.7) 31.7 (7.6) 50.5 (40.0) 65.6 (32.8) 7.1 (-) 72.4 (9.2) Fine-tuning 23.5 (9.2) 19.9 (3.6) 28.1 (5.0) 3.8 (0.8) 42.1 (6.7) 14.8 (2.3) 28.9 (10.9) 36.6 (7.4) 3.5 (1.3) 57.4 (17.7) EWC 28.3 (9.2) 19.5 (5.4) 25.5 (4.3) 6.9 (1.1) 39.8 (4.2) 18.3 (2.8) 18.6 (4.4) 24.6 (5.5) 5.5 (1.7) 52.1 (18.3) LiBOG 31.3 (11.9) 19.7 (5.8) 25.1 (5.8) 7.2 (1.0) 40.8 (3.9) 19.2 (1.6) 17.2 (4.5) 22.8 (5.5) 5.8 (1.7) 51.7 (14.8) LLR-BC 4.2 (1.0) 0.7 (0.5) 0.8 (0.4) 3.5 (0.8) 26.7 (4.5) 3.4 (1.3) 0.8 (0.3) 1.1 (0.4) 2.8 (1.6) 41.1 (21.0)
@@
Figure 3: Forgetting curve of task order 1, measured by average solution distance (vertical axis). Epochs 0–200 (first task) are omitted as no forgetting occurs. Restart is excluded due to significantly higher forgetting than other methods.
@@ 4.2 Stability and Plasticity
We analyze the stability by comparing the AF and AMF values over all five task orders, as shown in Table 1. LLR-BC achieves substantially lower AF and AMF values compared to all other methods, demonstrating its strong stability for effectively mitigating catastrophic forgetting. Figure 3 shows the average solution distance on seen tasks during lifelong learning of task order 1. At the start of each task, all baselines exhibit a sharp performance drop on previously learned tasks. In contrast, LLR-BC mitigates forgetting and maintains consistently lower distances with greater stability.
@@
Additionally, the rankings of methods differ on the two forgetting metrics, i.e., AF and AMF. For example, fine-tuning achieves a lower (better) AF than EWC but yields a higher (worse) AMF. This discrepancy arises because AF also accounts for backward transfer, i.e., training on a new task can sometimes improve performance on earlier, similar tasks. Backward transfer is also evident in Figure 3. This further underscores the importance of LLR-BC’s ability to maintain strong plasticity.
Similar patterns are observed across other task orders, as shown in Appendix D. In summary, LLR-BC consistently outperforms all baselines in both stability and plasticity.
@@ 4.3 Zero-shot Generalization
Beyond performance on learned tasks, we also expect neural solvers to acquire general knowledge from randomly ordered tasks, thereby improving their performance on unseen tasks. AG values for all methods are reported in Table 1. LLR-BC outperforms all baselines in terms of AG, demonstrating its strong ability to zero-shot generalize to a new task. We further test the solver after learning from task order 1 on TSPLIB and CVRPLIB, with the maximum scale as 1001. As the results presented in Table 2, LLR-BC performs the best on both benchmarks. Appendix D.3 presents the detailed values. Despite the uncontrollable and unpredictable nature of task order, LLR-BC demonstrates a strong generalization ability by capturing general knowledge across tasks.
Table 2: Mean (Std.) test performance on benchmark instances. All values are scaled by $10^{-3}$.
Benchmark Restart Fine-tuning EWC LiBOG LLR-BC
CVRPLIB 43.68 (27.37) 8.54 (10.71) 11.86 (11.59) 18.05 (10.86) 7.88 (8.75)
TSPLIB 99.76 (61.06) 38.16 (93.36) 27.80 (59.58) 27.96 (32.29) 18.08 (16.98)
@@ 4.4 Ablation Study
We conduct ablation studies on task order 1 to evaluate the contribution of the CaEW and DsBC. As shown in Table 3, removing either component leads to performance degradation across multiple metrics on both CVRP and TSP, confirming the effectiveness and contribution of both modules.
@@ 4.5 Hyperparameter Sensitivity Analysis
We varied the values of key hyperparameters, i.e., $|\mathcal{E}|$, $|\mathcal{B}|$ and $\alpha$, and evaluated the performance of LLR-BC across all metrics under each setting, as shown in Table 3. -nEW and -KLD denote the version without CaEW (using equal weights) and the version using KLD, respectively. -n(v) denotes the version with n set to v. Detailed discussion can be found in Appendix D.4. In summary, varying the hyperparameters leads to only minor performance fluctuations across evaluation metrics, compared with the performance gap between LLR-BC and baselines. This suggests that LLR-BC is robust and not overly sensitive to its hyperparameter settings.
@@
Method CVRP TSP Method CVRP TSP AP AF AMF APl AG AP AF AMF APl AG AP AF AMF APl AG AP AF AMF APl AG -$|\mathcal{E}|(4)$ 5.7 0.8 0.8 4.8 22.7 2.6 1.2 1.2 1.7 21.6 -$\alpha$(10) 13.2 5.1 5.5 4.2 34.7 1.3 0.4 0.6 1.2 20.5 -$|\mathcal{E}|(8)$ 5.8 1.0 1.1 4.2 22.8 2.0 0.9 1.1 1.5 21.6 -$\alpha$(50) 6.6 1.4 1.6 4.2 24.7 2.0 0.7 0.9 1.2 22.2 -$|\mathcal{B}|(250)$ 6.4 1.9 1.9 4.3 25.3 1.6 0.7 1.1 1.1 20.5 -$\alpha$(500) 5.1 0.2 0.3 4.9 22.5 3.5 1.3 1.3 3.2 21.2 -$|\mathcal{B}|(500)$ 5.8 1.3 1.3 4.2 29.8 2.2 0.6 0.9 1.7 22.2 -$\alpha$(1000) 5.9 0.3 0.4 5.8 23.6 4.9 1.4 1.4 4.6 22.3 -nEW 5.2 0.8 0.8 4.2 24.5 1.8 0.9 0.8 1.5 22.3 -KLD 5.5 0.7 0.7 4.3 23.1 1.9 0.9 0.9 1.5 22.3 Default 4.9 0.6 0.7 4.3 23.5 1.7 0.8 0.9 1.3 21.6

### v3

@@ 3.3 Decision-seeking Behavior Consolidation (DsBC)
Existing methods (Li et al., 2025; Feng et al., 2025) replay experiences by generating, solving, and learning instances of previous tasks, ignoring the instructive information contained in historical behaviors. In contrast, by minimizing the weighted sum of the difference between the current model’s behaviors and the buffered behavior at buffered states, LLR-BC encourages the consolidation of prior behavior patterns, thereby preserving knowledge and mitigating catastrophic forgetting.
The Kullback–Leibler divergence (KLD) $D_{KL}(P||Q)$, where $P$ and $Q$ are the teacher and leaner probability distribution respectively, is widely adopted for knowledge distillation of neural VRP solvers (Bi et al., 2022; Li et al., 2025; Zheng et al., 2025). However, recent studies (Kaplanis et al., 2019; Wang et al., 2024a; Qi et al., 2026) have shown that using reverse KLD (RKLD) $D_{RKL}(P||Q)=D_{KL}(Q||P)$ (detailed explanation is in Appendix A.3), could lead to better learning performance. While minimizing KLD pushes the learner distribution to spread probability mass across all modes of the given teacher distribution, minimizing RKLD leads to mode-seeking, where the learner tends to concentrate on the teacher’s highest-probability actions, while still maintaining relatively good alignment with the overall distribution (Chan et al., 2022; Wu et al., 2025). During problem-solving of constructive neural VRP solvers, the action (node) with the highest probability is typically chosen to extend the route. To address forgetting, that could manifest in significantly longer routes on previously seen tasks, it is crucial to preserve these highest-probability decisions. Therefore, instead of KLD, LLR-BC employs the RKLD to measure the behavioral difference in a decision-seeking way. To consolidate sampled behaviors, LLR-BC minimizes the following behavior consolidation loss term, considering the RKLD (Wang et al., 2024a; Qi et al., 2026):
@@
where $\{\tau\}$ is the set of newly obtained experience trajectories from the new task, $\mathcal{L}_{DRL}(\theta,\{\tau\})$ is the loss calculated by the adopted DRL algorithm based on $\{\tau\}$ to maximize the reward gain, and $\alpha$ is a hyperparameter to balance between consolidation of previous experiences and learning of the new task. Notably, LLR-BC is a generic framework and can be applied to different neural solver methods with different DRL algorithms, in which $\mathcal{L}_{DRL}$ in Eq. equation 2 can be implemented in different ways. Appendix A.5 detailedly discusses about computational cost of DsBC.
@@ 4 Experiments
  • Stability and plasticity: In lifelong learning process, how does LLR-BC perform in terms of stability and plasticity?
@@
To answer the questions, we simulate multiple lifelong learning scenarios on CVRP and TSP, and compare LLR-BC against widely used and representative lifelong learning baselines across diverse metrics [^1] Code will be released upon acceptance of this paper.. Most experiments are conducted based on POMO (Kwon et al., 2020), given its concise design and broad applicability. To verify the applicability of LLR-BC, we further evaluate LLR-BC on two representative and state-of-the-art constructive neural solvers designed for good cross-distribution cross-scale performance, i.e., Omni (Zhou et al., 2023) and INViT (Fang et al., 2024).
@@
Six distributions for sampling node coordinates are used: four proposed by Bossek et al. (2019) and widely used in neural VRP solver studies (Jiang et al., 2022; Zhou et al., 2023), i.e., Uniform (U), Gaussian Mixture (GM), Explosion (E), Compression (C), and two additionally designed ones for greater diversity: Grid (G) and Ring (R). For CVRP, node demand is also required to be sampled. While prior studies used uniformly random demand for all distributions (Kwon et al., 2020; Jiang et al., 2022; Zhou et al., 2023; Fang et al., 2024), we build six distinct demand distributions and assign them to the above six distributions to better reflect the diversity of real-world problems (Liu et al., 2021). Following common practice (Kwon et al., 2020; Jiang et al., 2022), we consider three problem scales: 20, 50, and 100. Specifically, tasks U and R correspond to scale 20, tasks G and E to scale 50, and tasks C and GM to scale 100, arbitrarily. We construct five task orders by randomly permuting the tasks. In addition, two classic and widely used benchmark datasets, TSPLIB (Reinelt, 1991) and CVRPLIB (Uchoa et al., 2017), are used for generalization ability evaluation. Appendix B provides the full definition of tasks and orders.
@@
Following Kwon et al. (2020), we train the solver on each task over 200 epochs, and solve a batch of sampled problems in parallel to leverage GPU parallelism during training. Therefore, both experience buffering and sampling operate at the batch level rather than the single experience level. Hyperparameters of LLR-BC are set as follows: $|\mathcal{B}|=1000$, $|\mathcal{E}|=16$, and $\alpha=100$, where $|\mathcal{B}|$ and $|\mathcal{E}|$ are defined in units of experience batches. The buffer constitutes only about 0.01% of the total training experiences obtained across all tasks. For evaluation, each task has a test set of 1,000 problem instances, which are different from the training ones. For each trial, all methods are evaluated on the same test sets. Additional details are provided in Appendix C.
@@
The goal of this work is not to design a totally new solver that outperforms the state-of-the-art solver on each individual task or a fixed set of tasks that can be learned simultaneously, but to improve the solver’s ability to learn unpredictable tasks sequentially. We consider several existing methods for comparison. First, we adapt the methods of (i) Li et al. (2025) (denoted Li (intra) and Li (inter)) and (ii) of Feng et al. (2025) (denoted Feng) to our scenarios and include them as baselines. Then, we involve several representative and commonly adopted strategies in neural solver studies (Kwon et al., 2020; Zhou et al., 2023): (iii) Restart, reinitialize the solver and train from scratch on each task; (iv) Fine-tuning, simply sequential training on each task. Additionally, we evaluate (v) EWC (Kirkpatrick et al., 2017), a widely used method in lifelong learning, especially in recent studies of optimization tasks (Manchanda and Ranu, 2023; Pei et al., 2025b). We also adapt and compare with (vi) LiBOG (Pei et al., 2025b), a recent method originally proposed for black-box optimization, within the VRP tasks. Appendix C.1 provides more details.
@@
  • Average Max Forgetting (AMF): the average of maximal forgetting of each task during the lifelong learning, i.e., $AMF=\frac{1}{k-1}\sum_{i=1}^{k-1}\max_{j=i+1}^{k}\max(0,\bar{d}_{j,i}-\bar{d}_{i,i})$.
@@
Smaller values of the above metrics indicate better performance. Unless otherwise specified, all reported values of the five metrics in tables are scaled by $10^{-3}$.
@@ 4.1 Performance in Solving Seen Tasks
The experimental results for CVRP and TSP are summarized in Table 1, with $k=6$ for all metrics. More detailed results with the absolute solution distance and optimality gap can be found in Appendices D.1, D.2. Across all task orders, LLR-BC achieves average AP values of $0.0042$ (CVRP) and $0.0034$ (TSP), whereas all other methods exceed $0.023$ (CVRP) and $0.014$ (TSP), respectively, indicating that LLR-BC obtains solutions of problems from learned tasks with substantially better quality. Moreover, LLR-BC exhibits greater robustness to task order variation. Compared methods show large fluctuations (larger Std.) in AP across orders, while LLR-BC consistently maintains low AP values with lower standard deviations. An additional comparison of LLR-BC with methods with non-lifelong settings, including POMO in multi-task training settings and INViT in its original training setting (Fang et al., 2024), further demonstrates LLR-BC’s superior performance (details can be found in Appendix D.3). These results demonstrate that LLR-BC effectively learns from sequentially arising tasks and retains the ability to solve all encountered tasks with high effectiveness.
@@
Method CVRP TSP AP AF AMF APl AG AP AF AMF APl AG Li (inter) 32.0 (6.8) 0.0 (0.0) 0.0 (0.0) 33.6 (7.7) 40.1 (10.5) 56.5 (38.3) 0.4 (0.5) 0.5 (0.5) 61.7 (45.6) 77.6 (59.6) Li (intra) 34.1 (2.1) 0.0 (0.0) 0.1 (0.1) 39.3 (1.7) 48.2 (4.8) 54.2 (7.2) 0.2 (0.3) 0.5 (0.4) 62.8 (12.5) 82.4 (27.7) Feng 24.6 (2.0) 3.2 (2.0) 4.5 (1.7) 24.2 (0.7) 39.1 (8.6) 24.1 (3.2) 1.8 (1.4) 3.8 (1.9) 21.4 (1.6) 45.5 (23.2) Restart 60.5 (29.3) 41.3 (26.3) 52.0 (25.5) 9.1 (-) 49.5 (4.7) 31.7 (7.6) 50.5 (40.0) 65.6 (32.8) 7.1 (-) 72.4 (9.2) Fine-tuning 23.5 (9.2) 19.9 (3.6) 28.1 (5.0) 3.8 (0.8) 42.1 (6.7) 14.8 (2.3) 28.9 (10.9) 36.6 (7.4) 3.5 (1.3) 57.4 (17.7) EWC 28.3 (9.2) 19.5 (5.4) 25.5 (4.3) 6.9 (1.1) 39.8 (4.2) 18.3 (2.8) 18.6 (4.4) 24.6 (5.5) 5.5 (1.7) 52.1 (18.3) LiBOG 31.3 (11.9) 19.7 (5.8) 25.1 (5.8) 7.2 (1.0) 40.8 (3.9) 19.2 (1.6) 17.2 (4.5) 22.8 (5.5) 5.8 (1.7) 51.7 (14.8) LLR-BC 4.2 (1.0) 0.7 (0.5) 0.8 (0.4) 3.5 (0.8) 26.7 (4.5) 3.4 (1.3) 0.8 (0.3) 1.1 (0.4) 2.8 (1.6) 41.1 (21.0)
@@
Figure 3: Forgetting curve of task order 1, measured by average solution distance (vertical axis). Epochs 0–200 (first task) are omitted as no forgetting occurs. Notably, some methods obtain too large solution distances and exceed the vertical range.
@@ 4.2 Stability and Plasticity
We analyze the stability by comparing the AF and AMF values over all five task orders, as shown in Table 1. Li (inter) and Li (intra) (Li et al., 2025), perform better in mitigating forgetting than LLR-BC. This is because half of their training resources (in terms of epochs) on the new task are used to learn purely from instances of previous tasks. Although they address forgetting well, they do not learn much from new tasks and perform poorly in terms of AP and APl. LLR-BC achieves substantially lower AF and AMF values than the rest of the compared methods, demonstrating its strong stability for effectively mitigating catastrophic forgetting. Figure 3 shows the average solution distance on seen tasks during lifelong learning of task order 1. At the start of each task, most baselines exhibit a sharp performance drop on previously learned tasks. Li (inter) and Li (intra) learn new tasks slowly. Feng suffers from low stability in maintaining learned knowledge. In contrast, LLR-BC mitigates forgetting and maintains consistently lower distances with greater stability.
@@
Additionally, the rankings of methods differ on the two forgetting metrics, i.e., AF and AMF. For example, fine-tuning achieves a lower (better) AF than EWC but yields a higher (worse) AMF. This discrepancy arises because AF also accounts for backward transfer, i.e., training on a new task can sometimes improve performance on earlier, similar tasks. Backward transfer is also evident in Figure 3. This further underscores the importance of LLR-BC’s ability to maintain strong plasticity. Similar patterns are observed across other task orders, as shown in Appendix D. In summary, LLR-BC consistently outperforms all baselines in both stability and plasticity.
@@ 4.3 Zero-shot Generalization
Beyond performance on learned tasks, we also expect neural solvers to acquire general knowledge from randomly ordered tasks, thereby improving their performance on unseen tasks. AG values for all methods are reported in Table 1. LLR-BC outperforms all baselines in terms of AG, demonstrating its strong ability to zero-shot generalize to a new task. We further test the solver after learning from task order 1 on TSPLIB and CVRPLIB, with the maximum scale as 1001. As shown by the results in Table 2, LLR-BC performs the best on both benchmarks. Appendix D.4 presents the detailed values. Despite the uncontrollable and unpredictable nature of task order, LLR-BC demonstrates a strong generalization ability by capturing general knowledge across tasks.
Table 2: Mean (Std.) test performance on benchmark instances.
Benchmark Restart Fine-tuning EWC LiBOG
CVRPLIB 43.68 (27.37) 8.54 (10.71) 11.86 (11.59) 18.05 (10.86)
TSPLIB 99.76 (61.06) 38.16 (93.36) 27.80 (59.58) 27.96 (32.29)
Benchmark Li (inter) Li (intra) Feng LLR-BC
CVRPLIB 39.31 (23.11) 46.80 (26.96) 27.38 (18.53) 7.88 (8.75)
TSPLIB 122.60 (51.78) 58.46 (26.55) 75.15 (30.83) 18.08 (16.98)
@@ 4.4 Ablation Study
We conduct ablation studies (denoted -nEW and -KLD) on task order 1 to evaluate the contribution of our key components, i.e. CaEW and DsBC. As shown in Table 3, removing either key component leads to performance degradation across multiple metrics on both CVRP and TSP, confirming the effectiveness and contribution of both modules. Furthermore, we conduct additional ablation studies to better investigate the characteristics of LLR-BC, including: (i) a variant that buffers experiences at every epoch instead of only the last epoch of each task (denoted -EE), (ii) variants that use entropy and top-2 margin instead of variance to quantify decision confidence (denoted -Ent and -T2M), (iii) a variant that rescales the reservoir sampling probabilities so that the buffered experiences are more uniformly distributed across tasks (denoted -Res), and (iv) a variant that buffers all steps for solve an instance as one experience, instead of just one step, denoted as -IB. The results show that both buffering only at the final epoch and using unscaled reservoir sampling contribute notably to the performance of LLR-BC. Replacing the confidence measure has only a minor impact, indicating that LLR-BC is not sensitive to the specific choice of confidence measurement. More details about ablation studies can be found in Appendix C.3.
@@ 4.5 Hyperparameter Sensitivity Analysis
We varied the values of key hyperparameters, i.e., $|\mathcal{E}|$, $|\mathcal{B}|$ and $\alpha$, and evaluated the performance of LLR-BC across all metrics under each setting, as shown in Table 3. -nEW and -KLD denote the version without CaEW (using equal weights) and the version using KLD, respectively. -n(v) denotes the version with n set to v. Detailed discussion can be found in Appendix D.5. In summary, varying the hyperparameters leads to only minor performance fluctuations across evaluation metrics, compared with the performance gap between LLR-BC and the baselines. This suggests that LLR-BC is robust and not overly sensitive to its hyperparameter settings.
@@
Method CVRP TSP Method CVRP TSP AP AF AMF APl AG AP AF AMF APl AG AP AF AMF APl AG AP AF AMF APl AG -$|\mathcal{E}|(4)$ 5.7 0.8 0.8 4.8 22.7 2.6 1.2 1.2 1.7 21.6 -$\alpha$(10) 13.2 5.1 5.5 4.2 34.7 1.3 0.4 0.6 1.2 20.5 -$|\mathcal{E}|(8)$ 5.8 1.0 1.1 4.2 22.8 2.0 0.9 1.1 1.5 21.6 -$\alpha$(50) 6.6 1.4 1.6 4.2 24.7 2.0 0.7 0.9 1.2 22.2 -$|\mathcal{B}|(250)$ 6.4 1.9 1.9 4.3 25.3 1.6 0.7 1.1 1.1 20.5 -$\alpha$(500) 5.1 0.2 0.3 4.9 22.5 3.5 1.3 1.3 3.2 21.2 -$|\mathcal{B}|(500)$ 5.8 1.3 1.3 4.2 29.8 2.2 0.6 0.9 1.7 22.2 -$\alpha$(1000) 5.9 0.3 0.4 5.8 23.6 4.9 1.4 1.4 4.6 22.3 -nEW 5.2 0.8 0.8 4.2 24.5 1.8 0.9 0.8 1.5 22.3 -KLD 5.5 0.7 0.7 4.3 23.1 1.9 0.9 0.9 1.5 22.3 -EE 7.8 3.1 3.1 3.9 24.1 2.6 2.1 2.4 1.6 22.7 -Ent 4.8 0.5 0.7 4.0 22.7 2.1 0.9 1.0 1.6 21.9 -Res 4.9 0.9 0.9 3.7 23.9 2.0 1.0 1.0 1.3 22.3 -T2M 4.8 1.0 1.2 3.4 22.8 1.7 0.9 1.1 1.0 21.6 Default 4.9 0.6 0.7 4.3 23.5 1.7 0.8 0.9 1.3 21.6 -IB 35.4 23.4 27.2 4.4 31.0 2.5 1.8 2.0 1.0 21.9